### v2

# Familiarity-Aware Evidence Compression for Retrieval-Augmented Generation

Abstract
Retrieval-augmented generation (RAG) improves large language models (LMs) by incorporating non-parametric knowledge through evidence retrieved from external sources. However, it often struggles to cope with inconsistent and irrelevant information that can distract the LM from its tasks, especially when multiple evidence pieces are required. While compressing the retrieved evidence with a compression model aims to address this issue, the compressed evidence may still be unfamiliar to the target model used for downstream tasks, potentially failing to utilize the evidence effectively. We propose FaviComp (Familiarity-aware Evidence Compression), a novel training-free evidence compression technique that makes retrieved evidence more familiar to the target model, while seamlessly integrating parametric knowledge from the model. Experimental results show that FaviComp consistently outperforms most recent evidence compression baselines across multiple open-domain QA datasets, improving accuracy by up to 28.1% while achieving high compression rates. Additionally, we demonstrate the effective integration of both parametric and non-parametric knowledge during evidence compression [^1] Code and data are available at https://github.com/luka-group/FaviComp.
Familiarity-Aware Evidence Compression for Retrieval-Augmented Generation
Dongwon Jung1  Qin Liu1  Tenghao Huang2  Ben Zhou3  Muhao Chen1 1University of California, Davis, 2University of Southern California, 3Arizona State University {dwojung,muhchen}@ucdavis.edutenghaoh@usc.edubenzhou@asu.edu
Figure 1: An overview of FaviComp. Instead of relying solely on compressed evidence from the compression model (upper), FaviComp familiarizes the compressed evidence to the target model while integrating parametric knowledge through ensemble decoding, resulting in improved downstream performance (lower).
1 Introduction
Retrieval-augmented generation (RAG) has become a common paradigm for large language models (LMs) to leverage external knowledge beyond their inherent knowledge boundaries to perform better in knowledge-intensive tasks such as open-domain question answering (QA) (Lewis et al., 2020; Izacard and Grave, 2021; Guu et al., 2020) and fact-checking (Pan et al., 2023; Li et al., 2024c). In particular, incorporating multiple evidence pieces is crucial in solving complicated tasks such as multi-hop and complex reasoning (Trivedi et al., 2023; Jiang et al., 2023b; Li et al., 2024b; Lu et al., 2023), which require various sources of information to solve the questions.
Nevertheless, RAG often struggles to cope with inconsistent and irrelevant information from the multiple evidence set, which can interfere with downstream tasks (Shi et al., 2023). This highlights the need for evidence compression to identify and retain only the essential information for LMs to utilize effectively. Traditionally, evidence compression has focused on reranking documents or sentences by relevance and then incorporating a top-ranked subset (Nogueira et al., 2020; Zhuang et al., 2023; Wang et al., 2023c) or compressing the documents into a compact form that retains only essential context (Jiang et al., 2023a; Xu et al., 2024; Yoon et al., 2024). However, the compressed evidence might be unfamiliar to the LM employed for the downstream task (referred to as the target model), particularly due to discrepancies in the internal knowledge and prompt preferences between the compression model and the target model (Gonen et al., 2023; Lee et al., 2024; Li et al., 2024a; Mallen et al., 2023). When LMs encounter unfamiliar contextual information, they often fail in balancing parametric and non-parametric knowledge, either by overly relying on their parametric knowledge (Longpre et al., 2021; Wang et al., 2023a; Zhou et al., 2023) or by utilizing retrieved evidence without considering its relevance to the input (Wu et al., 2024).
To address these challenges, we propose Familiarity-aware Evidence Compression (FaviComp), a training-free evidence compression method that consolidates multiple evidence into an abstractive summary that is more familiar to the target model, while seamlessly integrating parametric knowledge from the model. Inspired by the prior findings that an LM’s familiarity with a prompt is generally reflected by low perplexity (Liu et al., 2024; Gonen et al., 2023; Wang et al., 2023b), FaviComp proactively composes the compressed evidence in a way to lower the perplexity of the target model. Specifically, instead of directly selecting the highest probability token from the compression model at each decoding step, FaviComp selects the token from the ensemble of the token probabilities from both the compression and target models. This ensemble decoding therefore constrains the token search space of the compression model to those with lower perplexity for the target model, making the context more familiar to the target model (Liu et al., 2024).
Furthermore, FaviComp potentially synergizes the retrieved knowledge with the target model’s parametric knowledge introduced during ensemble decoding. It can effectively discern when to leverage internal or external knowledge, which is particularly beneficial in the presence of noisy contextual evidence in complex tasks such as multi-document or multi-hop QA (Wang et al., 2024).
Our experiments show that FaviComp outperforms most recent evidence compression baselines in five open-domain QA datasets, improving accuracy by up to 28.1% while maintaining high compression rates. Additionally, we conduct ablation studies by varying the degree of decoding ensemble and analyzing its impact on performance and context perplexity. Moreover, we investigate how FaviComp effectively integrates parametric and non-parametric knowledge during evidence compression.
2 Method
We present FaviComp, a decoding-time evidence compression method that familiarizes retrieved evidence with the target model while synergizing them with the model’s parametric knowledge. We first illustrate the motivation for FaviComp in Section 2.1 and provide the preliminaries of evidence compression in RAG Section 2.2, followed by a detailed definition of our proposed framework in Section 2.3.
2.1 Motivation and Method Overview
Figure 1 illustrates the overview of FaviComp. Existing evidence compression methods employ the compression model to filter out irrelevant information from the retrieved documents. However, since the compression model and the target model are different, the target model might not be familiar to the compressed evidence due to the difference in internal knowledge and prompt preferences between the two models (Gonen et al., 2023; Lee et al., 2024; Mallen et al., 2023). In addition, the compressed evidence cannot be supplemented with the rich parametric knowledge from the target model. In the example, even though the compression model successfully summarizes the essential information, the target model produces an inaccurate answer due to the unfamiliarity with the target model and the lack of integration of the parametric knowledge. On the other hand, FaviComp compresses the given evidence more favorable to the target model by using a novel ensemble decoding technique and leverages its parametric knowledge to supplement the missing evidence (“Lionel Messi made his league debut in Barcelona"), effectively combining evidential and parametric knowledge.
2.2 RAG with Evidence Compression
Given a set of $k$ retrieved evidence snippets $D=\{d_{1},d_{2},\ldots,d_{k}\}$ and a textual input sequence $x$, RAG aims to generate an output sequence $y$, conditioned on both $D$ and $x$. However, RAG directly utilizes $D$ which often contains irrelevant information to $x$, potentially confusing the target model in downstream tasks (Shi et al., 2023). Thus, we uses an additional compression model to condense $D$ into a concise and input-relevant context $c$, which is then used in place of $D$ during the downstream generation process. Thus, the RAG with evidence compression is formalized as:
$y^{*}=\arg\max_{y}P_{t}(y\mid x,\hat{c}),$
$\hat{c}=P_{c}(c\mid x,[d_{1},d_{2},\ldots,d_{k}]),$
where $y^{*}$ is the final output sequence, [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes concatenation, and $P_{t}$ and $P_{c}$ represent the probability distributions of the target and compression models, respectively. In this work, we consider any natural language prompting tasks, such as open-domain QA tasks, where $x$ represents the input prompt (also known as the query in QA tasks) and $y^{*}$ denotes the output sequence.
The compression model’s objective is to produce a concise yet informative summary $c$ of the evidential documents $D$ that captures the essential information relevant to the input query $x$. We use an unsupervised approach, where the model is instructed to generate a query-relevant summary of $D$ in a zero-shot manner using an evidence compression instruction prompt, denoted as $I_{comp}$, such as the one below:
Evidence Compression Instruction Given a question and multiple document snippets, generate one summarized context that is helpful to answer the question.
Specifically, the evidence compression is done in an auto-regressive way formalized as,
$P_{c}(c\mid I_{comp},x,D)=\prod_{i=1}^{|c|}P_{c}(c_{i}|I_{comp},x,D,c_{<i}),$
where $|c|$ is the length of the summary $c$.
2.3 Ensemble Decoding for FaviComp
Simple compression techniques might lead to subpar performance in downstream tasks because the compressed evidence may not be familiar to the target model. To better align the context to the target model, FaviComp proactively composes it to lower the target model’s perplexity by introducing a constraint in decoding space from the target model during the evidence compression. FaviComp achieves this goal through ensemble decoding, which involves a multiplicative ensemble of two LMs—compression model and target model—at each decoding step.
Specifically, the target model is directed to generate a context $c$ that would be helpful in answering the question $x$ without referencing the evidence set. This is also done in zero-shot using a context generation instruction prompt $I_{gen}$ such as:
Context Generation Instruction Given a question, generate a context that is helpful to answer the question.
The context generation is also performed in an auto-regressive fashion, represented as:
$P_{t}(c\mid I_{gen},x)=\prod_{i=1}^{|c|}P_{t}(c_{i}|I_{gen},x,c_{<i}),$
where $|c|$ denotes the length of the generated context $c$.
Once the compression model and the target model generate their respective probability distributions for the next token, the subsequent token is chosen by maximizing the weighted sum of the log probabilities from both models. The selected token is the continuation of the previously generated text aligned with their objectives. This process is formalized as follows:
$c_{i}$ $=\arg\max_{c_{i}\in V}(\alpha\cdot\log P_{t}(c_{i}|I_{gen},x,c_{<i})$
$+(1-\alpha)\cdot\log P_{c}(c_{i}|I_{comp},x,D,c_{<i})),$
where $c_{i}$ is the subsequent token, and $\alpha$ is the ensemble coefficient that weighs between the two probability distributions. We demonstrate how the coefficient $\alpha$ impacts both the perplexity and the downstream performance in Section 4.2.
Ensemble decoding proactively shifts the token search space in evidence compression by upweighting those tokens with lower perplexity from the target model’s perspective (Liu et al., 2024), resulting in a compressed evidence that is more familiar to the target model. Note that since both objectives ultimately share the goal of generating context relevant to the question, combining the logits ensures alignment with this ultimate goal.
In addition, ensemble decoding enables FaviComp to seamlessly integrate both retrieval knowledge from the external evidence set and the target model’s parametric knowledge. Specifically, FaviComp selects the $\arg\max$ token from the target model only when the token’s probability is higher than that of the compression model, demonstrating that FaviComp draws on parametric knowledge only when necessary—potentially when the compression model is uncertain about the next token. This is particularly beneficial for complex tasks like multi-document QA, where the evidence set may not include all the necessary information (Mallen et al., 2023). In such cases, the missing information in compressed evidence can be supplemented by tokens generated from context generation by the target model, which is entirely based on parametric knowledge. We demonstrate in Section 4.3 and Section 5 that FaviComp can incorporate knowledge from both sources effectively, leading to a performance boost compared to compression methods that solely focus on distilling knowledge from the evidence set.
3 Experimental Settings
We assess the effectiveness of FaviComp on knowledge-intensive QA tasks. In this section, we delve into the details of the experimental settings.
3.1 Datasets
We evaluate FaviComp on five open-domain QA datasets, including two single-document QA datasets, Natural Questions (NQ; Kwiatkowski et al. 2019) and TriviaQA (TQA; Joshi et al. 2017), and three multi-document QA datasets, HotpotQA (HQA; Yang et al. 2018), 2WikiMultiHopQA (Wiki; Ho et al. 2020), and MuSiQue (MQ; Trivedi et al. 2022). Following prior studies (Asai et al., 2023; Xu et al., 2024), we evaluate the performance on the development set of each dataset using two evaluation metrics, Accuracy (Acc) and token-level F1.
3.2 Implementation Details
For all the comparison methods, we utilize Llama3-8B-Instruct and Mixtral-8x7B- Instruct as the target model to tackle downstream QA tasks with RAG. For FaviComp and Zero-shot Summarization, we employ two compression models, one for each target model: Llama3.2-3B-Instruct for Llama3- 8B-Instruct target model and Mistral-7B- Instruct for Mixtral-8x7B-Instruct target model. For each question, we retrieve five documents from 2018 Wikipedia corpus (Karpukhin et al., 2020) using Contriever-MSMARCO (Izacard et al., 2021), so as to be consistent with previous studies (Xu et al., 2024; Yoon et al., 2024). We set ensemble coefficient $\alpha$ of FaviComp to 0.5 by default, for which more analyses are given in Section 4.2. The prompts used in the experiment are presented in Appendix C.
3.3 Baselines
We consider the following categories of baselines. (1) No Context: RAG without any context. (2) Gold Compression: RAG using directly relevant evidence from the retrieved documents if they exist. (3) Raw Document: RAG with raw documents that have not undergone any compression. (4) Generated Context (Yu et al., 2023): RAG with context generated by the same LM as the target model. This is equivalent to FaviComp with $\alpha=1$, as we rely solely on the target model to generate context when $\alpha=1$. (5) Reranking-based Methods: We rerank sentences in the evidence set and choose top-ranked sentences as the context. We utilize two rerankers—Sentence-BERT (Reimers and Gurevych, 2020) and RECOMP-extractive (Xu et al., 2024). (6) Compression-based Methods: We employ four compressors—LongLLMLingua (Jiang et al., 2023a), RECOMP-abstractive (Xu et al., 2024), CompAct (Yoon et al., 2024), and Zero-shot Summarization. For Zero-shot Summarization, we use the same evidence compression instruction prompt of FaviComp to summarize multiple evidence using the same LM as the target model. This is equivalent to FaviComp with $\alpha=0$, as we depend entirely on the compression model without any intervention from the target model. A detailed explanation of the implementation of the baselines is provided in Appendix A.
Methods Size NQ TQA HQA Wiki MQ
Acc F1 Acc F1 Acc F1 Acc F1 Acc F1
Llama3-8B-Instruct
Gold Compression - - - - - 42.3 51.3 35.7 40.0 10.2 17.7
No Context - 26.9 31.9 57.2 61.2 19.1 25.5 20.5 25.0 5.4 13.0
Raw Document - 42.6 47.1 67.6 70.8 30.3 38.7 22.0 26.8 8.2 15.0
Generated Context - 32.3 36.6 59.7 62.4 22.7 29.7 24.8 28.7 7.6 14.8
Sentence-BERT 110M 30.3 35.4 59.2 62.9 22.4 29.6 18.1 22.9 7.7 14.8
RECOMP-extractive 110M 33.7 38.1 59.4 62.8 22.5 29.8 18.0 22.4 8.1 15.5
LongLLMLingua 7B 35.4 40.9 64.8 67.6 25.9 34.7 19.2 24.2 7.7 14.4
RECOMP-abstractive 775M 39.3 43.3 62.9 66.1 27.0 34.8 20.5 25.0 7.3 14.8
CompAct 7B 42.3 46.1 67.0 69.7 29.8 37.5 21.4 26.6 9.2 16.9
Zero-shot Summarization 3B 39.4 43.2 64.2 67.1 30.1 38.5 25.7 31.1 7.7 15.3
FaviComp 3B 42.8 46.8 68.0 70.9 33.0 41.6 29.6 35.2 10.8 19.9
Mixtral-8x7B-Instruct
Gold Compression - - - - - 48.2 55.1 49.9 51.9 12.9 18.6
No Context - 36.7 38.4 68.9 72.0 25.1 31.6 32.5 35.9 6.4 11.8
Raw Document - 46.3 42.1 72.1 71.1 34.0 39.0 32.9 36.3 10.1 15.6
Generated Context - 33.6 33.9 61.4 62.9 26.5 32.9 30.2 34.3 7.2 13.4
Sentence-BERT 110M 36.8 36.8 67.0 68.7 28.3 34.5 32.5 36.2 9.9 15.2
RECOMP-extractive 110M 38.0 37.9 66.7 68.0 28.7 34.3 31.8 34.9 9.4 15.6
LongLLMLingua 7B 40.1 39.4 70.5 71.0 32.0 38.3 31.9 36.1 9.7 15.9
RECOMP-abstractive 775M 42.1 41.3 68.4 69.4 32.3 38.5 32.2 36.2 7.9 13.6
CompAct 7B 44.1 43.4 70.3 71.4 35.2 41.6 35.9 39.5 11.2 16.9
Zero-shot Summarization 7B 42.1 40.6 65.9 67.0 31.4 38.1 28.5 32.8 8.4 13.8
FaviComp 7B 43.6 44.5 72.6 73.9 36.3 44.4 40.5 45.2 13.4 19.9
Table 1: Experimental results on five open-domain QA datasets. Size column represents the size of the compression model used for each method. indicates a fully-supervised compression model, where the compressor is trained. The best Accuracy and token-level F1 scores for each dataset are in bold.
4 Experimental Results
In this section, we compare the overall performance of FaviComp with other baselines across the five datasets (Section 4.1), explore the impact of ensemble coefficient $\alpha$ on performance and perplexity (Section 4.2), investigate how effectively FaviComp incorporate parametric and non-parametric knowledge (Section 4.3), and compare the compression rates with other baselines (Section 4.4).
4.1 Main Results
The overall performance of FaviComp and the baselines across the five datasets are presented in Table 1.222We present additional experimental results using other combinations of compression and target model at Section B.1. To start with, the compression-based methods consistently outperform the reranking-based methods, due to the fact that the reranking-based methods are prone to losing more question-relevant information by discarding lower-ranked sentences.
Next, FaviComp outperforms all other baselines across all the datasets, except for the Gold Compression which is regarded as the upper bound of the performance. It is noteworthy that FaviComp, as a training-free strategy, outperforms all the supervised compression-based baselines that use similar or larger compression models [^3] We conduct a fair comparison with RECOMP-abstractive by using the same base compression model in Section B.2.. This result suggests that knowledge distillation from a larger teacher LM to a smaller compression model may not generalize well, as the context preferences and prior knowledge of the target model and the teacher model are likely to differ. In contrast, the superior performance of FaviComp is attributed to its ability to familiarize evidence with the target model and its effective incorporation of parametric knowledge from ensemble decoding. Moreover, for the MQ dataset, FaviComp even outperforms Gold Compression baseline which can be viewed as a perfect compressor. This demonstrates that explicitly incorporating parametric knowledge from the target model can significantly enhance performance in multi-document QA, even when the context is imperfect.
Finally, given that Zero-shot Summarization corresponds to FaviComp with $\alpha=0$ and Generated Context corresponds to FaviComp with $\alpha=1$, the fact that FaviComp outperforms both baselines highlights its ability to effectively incorporate tokens from both sources—evidence summary and generated context. This results in superior performance compared to relying on just one source alone.
Figure 2: Impact of coefficient $\alpha$ on performance and perplexity when using Llama3.2-3B-Instruct and Llama3-8B-Instruct compression-target pairs.
4.2 Impact of Ensemble Coefficient on Performance and Perplexity
Figure 2 illustrates how performance and perplexity change as the ensemble coefficient $\alpha$ is varied across the values when using Llama3.2-3B-Instruct and Llama3-8B- Instruct compression-target pairs on HQA and MQ datasets [^4] Results for other datasets are included in Figure 6.. We calculate the perplexity of the compressed evidence conditioned on the preceding inputs, i.e. instruction, demonstrations, and the question. For all the datasets, performance is the highest when $\alpha=0.5$, indicating that proactively lowering perplexity by equally weighting both input sources yields the best results. When $\alpha$ is below 0.5, performance improves as the perplexity of compressed evidence decreases, which aligns with the previous works (Liu et al., 2024; Gonen et al., 2023). However, when $\alpha$ exceeds 0.5, performance declines as perplexity decreases due to the lack of evidential knowledge during evidence compression. Additionally, when $\alpha$ reaches 0.9 or 1.0, there is a slight rise in the perplexity due to LM’s increased uncertainty with limited evidential knowledge.
Figure 3: Accuracy of baselines methods on $\mathrm{Hits=0}$ and $\mathrm{Hits=1}$ subset of multi-document QA datasets.
Methods NQ TQA HQA Wiki MQ LongLLMlingua 1.87 1.84 1.83 1.83 1.83 RECOMP-abstractive 17.96 17.79 19.72 32.06 32.05 CompAct 8.85 8.92 9.45 10.71 8.96 Zero-shot Summarization 14.12 17.12 18.75 21.39 16.19 FaviComp 16.43 22.40 22.55 23.10 18.95
Table 2: Compression rates of compression-based baselines. The highest compression rate is indicated in bold, while the second highest is underlined.
Question: This film is an adaption of a Jacques Offenbach’s opera that was written by a Hungarian British screenwriter?
Methods (Compressed) Evidence Prediction Perplexity
Raw Document …(skip)… The Tales of Hoffmann is a 1951 British Technicolor film adaptation of Jacques Offenbach’s opera “The Tales of Hoffmann", written, produced and directed by the team of Michael Powell and Emeric Pressburger working under the umbrella of their production company, The Archers. Emeric Pressburger 12.429
Zero-shot Summarization The 1951 film “The Tales of Hoffmann" is an adaptation of Jacques Offenbach’s opera, written, produced, and directed by Michael Powell and Emeric Pressburger. Emeric Pressburger 2.298
FaviComp The 1951 film “The Tales of Hoffmann" is an adaptation of Jacques Offenbach’s opera, written by Emeric Pressburger, a Hungarian-British screenwriter, and directed by Michael Powell and Emeric Pressburger. The Tales of Hoffmann 1.959
Question: Which magazine was first published earlier, The Chronicle of Philanthropy or Skeptic?
Raw Document The Chronicle of Philanthropy is a magazine that covers the nonprofit world. …(skip)… It was founded in 1988 by editor Phil Semas and then managing editor Stacy Palmer. …(skip)… Philanthropy (magazine) Philanthropy is a quarterly magazine published by the Philanthropy Roundtable. First published as a newsletter in 1987, "Philanthropy" became a glossy magazine in 1996. Philanthropy 4.856
Zero-shot Summarization The Chronicle of Philanthropy was founded in 1988, while Philanthropy magazine was first published as a newsletter in 1987 and became a glossy magazine in 1996. Philanthropy magazine 3.196
FaviComp The Chronicle of Philanthropy was first published in 1988, while Skeptic was first published in 1992. The Chronicle of Philanthropy 1.345
Table 3: Case study of evidence compression: FaviComp vs. Raw Document and Zero-shot Summarization. For FaviComp, the colors red and blue highlight tokens that are the $\arg\max$ of the compression model and the target model, respectively. Purple indicates a token that is the $\arg\max$ of neither model. Tokens with no coloring represent those that are the $\arg\max$ of both models.
4.3 Integration of Parametric and Non-parametric Knowledge
The effective integration of parametric and non-parametric knowledge is crucial for complex tasks such as multi-document QA, where the evidence set may not contain all the necessary information. To this end, we evaluate how effectively FaviComp incorporates parametric knowledge from the target model and non-parametric knowledge from the compression model on the multi-document QA datasets. We begin by dividing the test samples of each dataset into evidence-relevant and evidence-irrelevant subsets, using the $\mathrm{Hits}$ metric. The $\mathrm{Hits}$ metric is set to 1 (evidence-relevant) if the retrieved evidence set contains the correct answer, and 0 (evidence-irrelevant) if it does not. We then assess the downstream performance of each subset. The underlying intuition is that if a method performs better on the evidence-relevant subset, it suggests that the method is more effectively utilizing the provided evidential knowledge. Conversely, if a method excels on the evidence-irrelevant subset, it indicates that the method is more effectively leveraging parametric knowledge without relying on potentially irrelevant evidence.
As shown in Figure 3, we compare the accuracy of FaviComp with Llama3.2-3B-Instruct and Llama3-8B-Instruct compression-target pairs on $\mathrm{Hits=0}$ and $\mathrm{Hits=1}$ subsets with the top-performing baselines, Zero-shot Summarization and CompAct [^5] We provide results of FaviComp on vaious alpha values in Section B.3. FaviComp outperforms other baselines in the $\mathrm{Hits=0}$ subset while performing comparably with others in the $\mathrm{Hits=1}$ subset. This proves that FaviComp effectively relies on parametric knowledge rather than evidential knowledge when faced with irrelevant evidence, while maintaining similar effectiveness in utilizing evidential knowledge when relevant evidence is present.
Interestingly, even though CompAct generally performs better on the $\mathrm{Hits=1}$ subset compared to Zero-shot Summarization, it underperforms relative to Zero-shot Summarization on the $\mathrm{Hits=0}$ subset. This suggests that the training may have been biased towards utilizing solely evidential knowledge, rather than effectively leveraging both sources in synergy.
4.4 Compression Rate Comparisons
[…]
5 Case Study
Table 3 presents two examples from HQA to illustrate how FaviComp effectively familiarizes evidence while seamlessly integrating both parametric and non-parametric knowledge during evidence compression. We compare its output with Raw Document, which does not apply any compression, and Zero-shot Summarization.
In both examples, Raw Document fails to produce the correct answer, even though the evidence contains the necessary information, highlighting the need for effective evidence compression. In the first example, while the difference between the compressed evidence from Zero-shot Summarization and FaviComp appears subtle, FaviComp delivers the correct answer with a lower perplexity in compression, underscoring the significance of evidence familiarization. The second example highlights the importance of parametric knowledge when the retrieved evidence set lacks complete information. Since the evidence set does not mention "Skeptic", Zero-shot Summarization introduces irrelevant information ("Philanthropy magazine"), ultimately leading to an incorrect answer. In contrast, FaviComp integrates parametric knowledge about "Skeptic" and incorporates it into the evidence compression. Notably, FaviComp selects the $\arg\max$ token from the target model only when the token’s probability is higher than that of the compression model, demonstrating that FaviComp draws on parametric knowledge only when necessary—potentially when the compression model is uncertain about the next token.
6 Related Works
Evidence Compression for RAG. Recent efforts on evidence compression seek to compress retrieved evidence pieces to filter out unnecessary information and retain only the essential context Wang et al. (2023c); Li et al. (2024d); Ke et al. (2024); Xu et al. (2024); Yoon et al. (2024). Most recently, Jiang et al. (2023a) and Cao et al. (2024) conduct token-level or embedding-based compression to preserve only the query-relevant information using a trained compressor. Xu et al. (2024) and Yoon et al. (2024) train a compression model to generate an abstractive summary of the documents by distilling knowledge from larger language models.
While these methods are successful to some extent, they often achieve suboptimal performance because the compressed context may be unfamiliar to the LM used in the downstream task due to differences in the internal knowledge and prompt preferences between the compression and the target model. In contrast, FaviComp proactively compresses the evidence pieces in a way to lower the target model’s perplexity using an ensemble decoding technique without any training, thereby improving the downstream performance.
Parametric and Non-parametric Knowledge in RAG. While there have been studies on the phenomena of LM’s utilization of both parametric and non-parametric knowledge sources, there is a lack of research focused on effectively synergizing both sources. A few of these efforts introduce counterfactual augmentation (Longpre et al., 2021; Fang et al., 2024; Zhang et al., 2024) and causal intervention (Zhou et al., 2023; Wang et al., 2023a) to mitigate knowledge conflict, which, however, requires explicitly knowing the features of the input that causes such conflict. Zhang et al. (2023) seek to address this issue by incorporating LM-generated context into the LM’s input along with the retrieved documents, thereby integrating both sources of knowledge. However, merely concatenating both contexts is a suboptimal solution, as LMs may still show bias toward one source over the other when generating responses (Longpre et al., 2021; Wu et al., 2024). To address this, FaviComp employs ensemble decoding during the evidence compression, ensuring that both types of knowledge are seamlessly fused together to create a consistent context.
Constrained Decoding. Constrained decoding has been previously proposed in text generation tasks for various purposes, including optimizing prompts (Liu et al., 2024), enhancing plausibility (Li et al., 2023) or controllability (Meng et al., 2022; Huang et al., 2023), and reducing hallucination (Shi et al., 2024). Our work is closely connected with the method by Liu et al. (2024) which employs ensemble decoding to paraphrase prompts to enhance zero-shot LM prompting and generalization. Their approach focuses on the robustness and generalizability of instruction prompts for tasks without retrieval augmentation. In contrast, our approach compresses externally retrieved evidence while integrating parametric knowledge during compression, specifically targeting knowledge-intensive tasks that require balancing both evidential and parametric knowledge.
7 Conclusion
In this study, we introduce FaviComp, a training-free, decoding-time evidence compression method designed to enhance RAG performance by consolidating retrieved evidence set to be more familiar to the target model, while seamlessly integrating parametric knowledge. Our extensive experiments validate the effectiveness of FaviComp on open-domain QA tasks, showing significant improvements over recent evidence compression baselines in multiple datasets. Additionally, FaviComp’s model-agnostic nature allows it to be effortlessly incorporated into various RAG workflows without additional training, making it a versatile tool for enhancing LMs in complex tasks.
Limitations
Although FaviComp exhibits superior performance in RAG compared to the recent evidence compression baselines, it has some limitations. (1) FaviComp consumes approximately twice as much computation compared to methods that only use a compression model since it needs two inferences (compression and target model) during the ensemble decoding. However, it is a training-free strategy that can be easily plugged into any RAG application. (2) Ensemble decoding requires the compression and target model to share the same vocabulary and tokenizer, which can limit the range of compatible models. Nonetheless, recent studies, such as Gu et al. (2024), have introduced techniques to enable model-agnostic ensemble decoding. This implies that there will be a potential direction of incorporating model-agnostic ensemble decoding with our framework to enable more flexible integration of various models, which we leave as future work.
Ethics Statement
This work follows the ACL Code of Ethics. We believe no potential risk is directly associated with the presented work.
Appendix A Implementation Details
(1) Gold Compression: We implement the Gold Compression baseline following the approach outlined by Yoon et al. (2024). We evaluate only on HQA, Wiki, and MQ, as these datasets contain gold documents. We first identify the presence of any gold documents in the retrieved documents. If found, we use the documents as the context. If none of the retrieved documents are identified as gold, we utilize the entire set of retrieved documents as the context for the evaluation. To identify the gold documents within the retrieved documents, we compare each gold document with the retrieved ones. If 50% or more of the content matches, we classify it as a gold document. This approach is necessary because the documents are chunked, and the retrieved documents may not exactly match the gold documents.
(2) Generated Context: We use the context generation prompt in Table 7 to generate the context.
(3) Zero-shot Summarization: We use the evidence compression prompt in Table 7 to compress the retrieved documents.
(4) RECOMP-extractive: We utilize the same Contriever models trained by the authors for each dataset, to encode both the question and the sentences in the evidence set. For Wiki and MQ, since there are no fine-tuned models available, we use the Contriever fine-tuned on HQA. Following the original paper, we select one sentence as the context for NQ and TQA, whereas for the other datasets, we utilize two sentences.
(5) RECOMP-abstractive: Similar to RECOMP-extractive, we use the same T5-large models trained by the authors for each dataset to compress the retrieved evidence. For the Wiki and MQ, we employ the T5-large model fine-tuned on HQA.
(6) LongLLMLingua: We use Llama2-7B [^6] https://huggingface.co/NousResearch/Llama-2-7b-hf trained by the authors as the prompt compressor model. We use the default hyperparameters in the original paper, where the dynamic context compression rate is set to 0.3, and the maximum compression rate is set to 0.5.
(7) CompAct: We use the same Mistral-7B- Instruct [^7] https://huggingface.co/cwyoon99/CompAct-7b model instruction-tuned by the authors for evidence compression. The number of documents per segment is set to 5 with 1 iteration.
Appendix B Additional Experiment Results
B.1 Other Compression and Target Models
We conduct an experiment where we use Llama3 -8B-Instruct and Mistral-7B-Instruct for both compression and target models. The result in Table 5 demonstrates that FaviComp outperforms all other baselines, supplementing the effectiveness shown in Section 4.1.
B.2 Head-to-Head Comparison with RECOMP-abstractive
Since the lower performance of RECOMP-abstractive might possibly be due to the use of smaller base model for compression (T5-large), we conduct a head-to-head experiment on FaviComp and RECOMP-abstractive by using the same base compression model. We construct training data on NQ, TQA, and HQA according to Xu et al. (2024) and finetune Mistral-7B-Instruct on each of the training data. We train for 7 epochs using LoRA with Adam optimizer with a learning rate of 2e-6 and a batch size of 64. We present the evaluation results in Table 4. Even though using larger base model for compression enhances the performance of RECOMP-abstractive to some extent, it still underperforms compared to training-free FaviComp. This underscores that the familiarization during evidence compression and integration of parametric and non-parametric knowledge are more helpful to the downstream generation than relying on a trained model for evidence compression.
B.3 Performance of $\mathrm{Hits=0}$ and $\mathrm{Hits=1}$ on Varying Alpha Values
We evaluate FaviComp’s performance on evidence-relevant ($\mathrm{Hits=1}$) and evidence-irrelevant ($\mathrm{Hits=0}$) subsets by varying $\alpha$ values. Figure 4 shows that $\alpha=0.5$ or $\alpha=0.7$ performs the best on the $\mathrm{Hits=0}$ subset, while performance declines as $\alpha$ deviates further from the value. This pattern in the $\mathrm{Hits=0}$ subset mirrors the overall performance trend, suggesting that appropriately utilizing parametric knowledge when the evidence is irrelevant is crucial to the overall performance. In the $\mathrm{Hits=1}$ subset, performance remains consistent for $\alpha$ values up to 0.5 but decreases significantly when $\alpha$ exceeds 0.5 due to the diminished utilization of the relevant evidential context.
Figure 4: Accuracy of FaviComp with various $\alpha$ values on $\mathrm{Hits=0}$ and $\mathrm{Hits=1}$ subset of multi-document QA datasets.
Appendix C Prompt Templates
Evaluation Prompt Template {System Prompt} {Demonstrations} Question: {Question} Context: {Context} Answer:
Figure 5: Evaluation Prompt Template.
C.1 Evaluation
The evaluation prompt template is shown in Figure 5. For all the evaluations throughout the experiment, we switch the positions of the Question and Context if doing so results in better performance. System prompts and demonstrations used in the evaluations are presented in Table 6 and Table 8, respectively.
C.2 FaviComp
The prompt templates for evidence compression and context generation of FaviComp are presented in Table 7
Appendix D Licenses
We include the licenses of datasets and models we used in this work.
Dataset Licenses:
  • NQ: Apache-2.0
  • TQA: Apache-2.0
  • HQA: CC BY-SA 4.0
  • Wiki: Apache-2.0
  • MQ: CC-BY-4.0
Model Licenses:
Methods NQ TQA HQA Wiki MQ
Acc F1 Acc F1 Acc F1 Acc F1 Acc F1
Llama3-8B-Instruct
Gold Compression - - - - 42.3 51.3 35.7 40.0 10.2 17.7
No Context 26.9 31.9 57.2 61.2 19.1 25.5 20.5 25.0 5.4 13.0
Raw Document 42.6 47.1 67.6 70.8 30.3 38.7 22.0 26.8 8.2 15.0
Generated Context 32.3 36.6 59.7 62.4 22.7 29.7 24.8 28.7 7.6 14.8
Sentence-BERT 30.3 35.4 59.2 62.9 22.4 29.6 18.1 22.9 7.7 14.8
RECOMP-extractive 33.7 38.1 59.4 62.8 22.5 29.8 18.0 22.4 8.1 15.5
LongLLMLingua 35.4 40.9 64.8 67.6 25.9 34.7 19.2 24.2 7.7 14.4
RECOMP-abstractive 39.3 43.3 62.9 66.1 27.0 34.8 20.5 25.0 7.3 14.8
CompAct 42.3 46.1 67.0 69.7 29.8 37.5 21.4 26.6 9.2 16.9
Zero-shot Summarization 41.3 45.1 66.3 69.5 30.2 38.6 22.3 28.1 8.3 16.3
FaviComp 42.3 46.6 68.4 71.5 32.3 41.0 27.6 33.6 11.4 20.1
Mistral-7B-Instruct
Gold Document - - - - 41.0 50.5 38.1 40.3 9.6 15.2
No Context 28.1 27.5 58.8 60.9 19.7 24.8 21.9 22.8 5.2 9.7
Raw Document 40.2 39.3 66.2 68.6 30.3 37.2 26.6 28.5 7.5 13.1
Generated Context 30.1 31.7 57.3 60.7 23.7 30.6 25.1 29.5 7.1 12.8
Sentence-BERT 29.8 30.1 57.8 60.7 23.8 30.3 22.9 24.7 7.5 12.3
RECOMP-extractive 31.7 32.2 57.2 60.0 24.1 30.2 23.2 24.4 7.4 12.5
LongLLMLingua 34.3 36.4 63.8 66.9 27.0 34.7 25.5 28.0 7.1 13.0
RECOMP-abstractive 38.0 37.8 62.1 65.0 27.4 34.3 25.1 27.4 6.4 12.0
CompAct 38.8 38.9 65.1 67.1 30.2 37.1 24.9 27.6 8.2 13.6
Zero-shot Summarization 38.4 38.2 62.3 64.8 28.2 35.2 23.2 27.1 6.8 11.8
FaviComp 40.3 40.4 65.9 68.9 32.0 40.5 29.7 35.1 9.2 15.2
Table 5: Additional experimental results. Llama3-8B-Instruct and Mistral-7B-Instruct are used for both compression and target models.
Figure 6: Impact of coefficient $\alpha$ on performance and perplexity for NQ, TQA and Wiki.
Target Models System Prompt
Llama-3-8B-Instruct You are an expert in Question Answering. Your job is to answer questions in 1 to 5 words based on the given context.
Mixtral-8x7B-Instruct You are an expert in Question Answering. Your job is to answer questions in 1 to 5 words based on the given context. Just output the answer as concisely as possible, no other words
Mistral-7B-Instruct You are an expert in Question Answering. Your job is to answer questions in 1 to 5 words based on the given context. Just output the answer as concisely as possible, no other words
Table 6: System prompts used in evaluation
Instruction Prompt Template
Evidence Compression You are an expert in summarization. Given a question and multiple document snippets, generate one summarized context that is helpful to answer the question. Just summarize, no other words. Question: {Question} Documents: {Evidence} Summarized Context:
Context Generation You are an expert in context generation. Given a question, generate a context that is helpful to answer the question. Just generate the context, no other words. Question: {Question} Context:
Table 7: Prompt Templates for FaviComp
Dataset Demonstrations
NQ Question: who sings i’ve got to be me Answer: Sammy Davis, Jr Question: who wrote i will follow you into the dark Answer: Ben Gibbard Question: who won season 2 of total drama island Answer: Owen (Scott McCord) Question: what part of the mammary gland produces milk Answer: cuboidal cells Question: when did the golden compass book come out Answer: 1995
TQA Question: Who sang the theme for the James Bond film ‘Thunderball’? Answer: Tom Jones Question: A hendecagon has how many sides? Answer: Eleven Question: In the 1968 feature film Chitty Chitty Bang Bang, of what country is Baron Bomburst the tyrant ruler? Answer: Vulgaria Question: Artists Chuck Close, Henri-Edmond Cross, John Roy, Georges-Pierre Seurat, Paul Signac, Maximilien Luce and Vincent van Gogh painted in what style? Answer: Pointillism Question: What is the study of the relation between the motion of a body and the forces acting on it? Answer: Dynamics
HQA Question: Which magazine was started first Arthur’s Magazine or First for Women? Answer: Arthur’s Magazine Question: The Oberoi family is part of a hotel company that has a head office in what city? Answer: Delhi Question: Musician and satirist Allie Goertz wrote a song about the "The Simpsons" character Milhouse, who Matt Groening named after who? Answer: President Richard Nixon Question: Are Jane and First for Women both women’s magazines? Answer: Yes Question: Were Pavel Urysohn and Leonid Levin known for the same type of work? Answer: No
Wiki Question: Where was the place of death of Marie Thérèse Of France (1667–1672)’s father? Answer: Palace of Versailles Question: Who is the paternal grandmother of Przemysław Potocki? Answer: Ludwika Lubomirska Question: Who lived longer, Herbert Findeisen or Léonie Humbert-Vignot? Answer: Léonie Humbert-Vignot Question: Are Alison Skipper and Diane Gilliam Fisher from the same country? Answer: Yes Question: Are director of film Move (1970 Film) and director of film Méditerranée (1963 Film) from the same country? Answer: No
MQ Question: Who is the child of the director and star of Awwal Number? Answer: Suneil Anand Question: What county shares a border with the county where Black Hawk Township is located? Answer: Dodge County Question: Who is the sibling of the person credited with the reinvention and popularization of oil paints? Answer: Hubert Van Eyck Question: Who heads the Catholic Church, in the country that a harp is associated with, as a lion is associated with the country that Queen Margaret and her son traveled to? Answer: Eamon Martin
Table 8: Demonstrations used in evaluation for each dataset

### v3

@@
 While compressing the retrieved evidence with a compression model aims to address this issue, the compressed evidence may still be unfamiliar to the target model used for downstream tasks, potentially failing to utilize the evidence effectively.
-We propose FaviComp (Familiarity-aware Evidence Compression), a novel training-free evidence compression technique that makes retrieved evidence more familiar to the target model, while seamlessly integrating parametric knowledge from the model.
-Experimental results show that FaviComp consistently outperforms most recent evidence compression baselines across multiple open-domain QA datasets, improving accuracy by up to 28.1% while achieving high compression rates.
-Additionally, we demonstrate the effective integration of both parametric and non-parametric knowledge during evidence compression [^1] Code and data are available at https://github.com/luka-group/FaviComp.
+We propose FaviComp (Familiarity-aware Evidence Compression), a novel inference-time evidence compression technique that makes retrieved evidence more familiar to the target model, while seamlessly integrating parametric knowledge from the model.
+Experimental results show that FaviComp consistently outperforms the most recent evidence compression baselines across multiple open-domain QA datasets, improving accuracy by up to 28.1% while achieving high compression rates.
+Additionally, we demonstrate the effective integration of both parametric and non-parametric knowledge during evidence compression.
+[^1] Code and data are available at https://github.com/luka-group/FaviComp
 Familiarity-Aware Evidence Compression for Retrieval-Augmented Generation
-Dongwon Jung1  Qin Liu1  Tenghao Huang2  Ben Zhou3  Muhao Chen1 1University of California, Davis, 2University of Southern California, 3Arizona State University {dwojung,muhchen}@ucdavis.edutenghaoh@usc.edubenzhou@asu.edu
+Dongwon Jung1   Qin Liu1   Tenghao Huang2   Ben Zhou3   Muhao Chen1 1University of California, Davis, 2University of Southern California, 3Arizona State University {dwojung,qinli,muhchen}@ucdavis.edutenghaoh@usc.edubenzhou@asu.edu
 Figure 1: An overview of FaviComp.
@@
 In particular, incorporating multiple evidence pieces is crucial in solving complicated tasks such as multi-hop and complex reasoning (Trivedi et al., 2023; Jiang et al., 2023b; Li et al., 2024b; Lu et al., 2023), which require various sources of information to solve the questions.
-Nevertheless, RAG often struggles to cope with inconsistent and irrelevant information from the multiple evidence set, which can interfere with downstream tasks (Shi et al., 2023).
+Nevertheless, RAG often struggles to cope with inconsistent and irrelevant information from the multiple evidence pieces, which can interfere with downstream tasks (Shi et al., 2023).
 This highlights the need for evidence compression to identify and retain only the essential information for LMs to utilize effectively.
@@
 When LMs encounter unfamiliar contextual information, they often fail in balancing parametric and non-parametric knowledge, either by overly relying on their parametric knowledge (Longpre et al., 2021; Wang et al., 2023a; Zhou et al., 2023) or by utilizing retrieved evidence without considering its relevance to the input (Wu et al., 2024).
-To address these challenges, we propose Familiarity-aware Evidence Compression (FaviComp), a training-free evidence compression method that consolidates multiple evidence into an abstractive summary that is more familiar to the target model, while seamlessly integrating parametric knowledge from the model.
+To address these challenges, we propose Familiarity-aware Evidence Compression (FaviComp), an inference-time evidence compression method that consolidates multiple evidence into an abstractive summary that is more familiar to the target model, while seamlessly integrating parametric knowledge from the model.
 Inspired by the prior findings that an LM’s familiarity with a prompt is generally reflected by low perplexity (Liu et al., 2024; Gonen et al., 2023; Wang et al., 2023b), FaviComp proactively composes the compressed evidence in a way to lower the perplexity of the target model.
@@
 2 Method
-We present FaviComp, a decoding-time evidence compression method that familiarizes retrieved evidence with the target model while synergizing them with the model’s parametric knowledge.
-We first illustrate the motivation for FaviComp in Section 2.1 and provide the preliminaries of evidence compression in RAG Section 2.2, followed by a detailed definition of our proposed framework in Section 2.3.
+We present FaviComp, a inference-time evidence compression method that familiarizes retrieved evidence with the target model while synergizing them with the model’s parametric knowledge.
+We first illustrate the motivation for FaviComp in Section˜2.1 and provide the preliminaries of evidence compression in RAG Section˜2.2, followed by a detailed definition of our proposed framework in Section˜2.3.
 2.1 Motivation and Method Overview
-Figure 1 illustrates the overview of FaviComp.
+Figure˜1 illustrates the overview of FaviComp.
 Existing evidence compression methods employ the compression model to filter out irrelevant information from the retrieved documents.
@@
 However, RAG directly utilizes $D$ which often contains irrelevant information to $x$, potentially confusing the target model in downstream tasks (Shi et al., 2023).
-Thus, we uses an additional compression model to condense $D$ into a concise and input-relevant context $c$, which is then used in place of $D$ during the downstream generation process.
+Thus, we use an additional compression model to condense $D$ into a concise and input-relevant context $c$, which is then used in place of $D$ during the downstream generation process.
 Thus, the RAG with evidence compression is formalized as:
-$y^{*}=\arg\max_{y}P_{t}(y\mid x,\hat{c}),$
-$\hat{c}=P_{c}(c\mid x,[d_{1},d_{2},\ldots,d_{k}]),$
-where $y^{*}$ is the final output sequence, [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes concatenation, and $P_{t}$ and $P_{c}$ represent the probability distributions of the target and compression models, respectively.
+[…]
+$\hat{c}=P_{\text{comp}}(c\mid x,[d_{1},d_{2},\ldots,d_{k}]),$
+where $y^{*}$ is the final output sequence, $\cdot$$\cdot$] denotes concatenation, and $P_{\text{tar}}$ and $P_{\text{comp}}$ represent the probability distributions of the target and compression models, respectively.
 In this work, we consider any natural language prompting tasks, such as open-domain QA tasks, where $x$ represents the input prompt (also known as the query in QA tasks) and $y^{*}$ denotes the output sequence.
@@
 Specifically, the evidence compression is done in an auto-regressive way formalized as,
-$P_{c}(c\mid I_{comp},x,D)=\prod_{i=1}^{|c|}P_{c}(c_{i}|I_{comp},x,D,c_{<i}),$
-where $|c|$ is the length of the summary $c$.
+$P_{\text{comp}}(c\mid\mathcal{C}_{\text{comp}})=\prod_{i=1}^{|c|}P_{\text{comp}}(c_{i}\mid\mathcal{C}_{\text{comp}},c_{<i}),$
+where $\mathcal{C}$ denotes the input prompt, constructed by stringifying $\{I_{\text{comp}},x,D\}$ using a predefined prompt template and $|c|$ is the length of the summary $c$.
 2.3 Ensemble Decoding for FaviComp
@@
 FaviComp achieves this goal through ensemble decoding, which involves a multiplicative ensemble of two LMs—compression model and target model—at each decoding step.
-Specifically, the target model is directed to generate a context $c$ that would be helpful in answering the question $x$ without referencing the evidence set.
+Specifically, the target model is instructed to generate a context $c$ that would be helpful in answering the question $x$ without referencing the evidence set.
 This is also done in zero-shot using a context generation instruction prompt $I_{gen}$ such as:
@@
 The context generation is also performed in an auto-regressive fashion, represented as:
-$P_{t}(c\mid I_{gen},x)=\prod_{i=1}^{|c|}P_{t}(c_{i}|I_{gen},x,c_{<i}),$
-where $|c|$ denotes the length of the generated context $c$.
+$P_{\text{tar}}(c\mid\mathcal{C}_{\text{gen}})=\prod_{i=1}^{|c|}P_{\text{tar}}(c_{i}|\mathcal{C}_{\text{gen}},c_{<i}),$
+where $\mathcal{C}_{\text{gen}}$ denotes the input prompt constructed using $\{I_{\text{gen}},x\}$ [^2] We provide the prompt templates for evidence compression and context generation in Table 10.
+and $|c|$ denotes the length of the generated context $c$.
 Once the compression model and the target model generate their respective probability distributions for the next token, the subsequent token is chosen by maximizing the weighted sum of the log probabilities from both models.
@@
 This process is formalized as follows:
-$c_{i}$ $=\arg\max_{c_{i}\in V}(\alpha\cdot\log P_{t}(c_{i}|I_{gen},x,c_{<i})$
-$+(1-\alpha)\cdot\log P_{c}(c_{i}|I_{comp},x,D,c_{<i})),$
+$c_{i}$ $=\arg\max_{c^{\prime}_{i},c^{\prime\prime}_{i}\in V}(\alpha\cdot\log P_{\text{tar}}(c^{\prime}_{i}\mid\mathcal{C}_{\text{gen}},c_{<i})$
+$+(1-\alpha)\cdot\log P_{\text{comp}}(c^{\prime\prime}_{i}\mid\mathcal{C}_{\text{comp}},c_{<i})),$
 where $c_{i}$ is the subsequent token, and $\alpha$ is the ensemble coefficient that weighs between the two probability distributions.
-We demonstrate how the coefficient $\alpha$ impacts both the perplexity and the downstream performance in Section 4.2.
-Ensemble decoding proactively shifts the token search space in evidence compression by upweighting those tokens with lower perplexity from the target model’s perspective (Liu et al., 2024), resulting in a compressed evidence that is more familiar to the target model.
+We demonstrate how the coefficient $\alpha$ impacts both the perplexity and the downstream performance in Section˜4.2.
+Ensemble decoding proactively shifts the token search space in evidence compression by upweighting those tokens with lower perplexity from the target model’s perspective, resulting in a compressed evidence that is more familiar to the target model.
 Note that since both objectives ultimately share the goal of generating context relevant to the question, combining the logits ensures alignment with this ultimate goal.
@@
 In such cases, the missing information in compressed evidence can be supplemented by tokens generated from context generation by the target model, which is entirely based on parametric knowledge.
-We demonstrate in Section 4.3 and Section 5 that FaviComp can incorporate knowledge from both sources effectively, leading to a performance boost compared to compression methods that solely focus on distilling knowledge from the evidence set.
+We demonstrate in Section˜4.3 and Section˜5 that FaviComp can incorporate knowledge from both sources effectively, leading to a performance boost compared to compression methods that solely focus on distilling knowledge from the evidence set.
 3 Experimental Settings
@@
 For each question, we retrieve five documents from 2018 Wikipedia corpus (Karpukhin et al., 2020) using Contriever-MSMARCO (Izacard et al., 2021), so as to be consistent with previous studies (Xu et al., 2024; Yoon et al., 2024).
-We set ensemble coefficient $\alpha$ of FaviComp to 0.5 by default, for which more analyses are given in Section 4.2.
-The prompts used in the experiment are presented in Appendix C.
+We set ensemble coefficient $\alpha$ of FaviComp to 0.5 by default, for which more analyses are given in Section˜4.2.
+The prompts used in the experiment are presented in Appendix˜C.
 3.3 Baselines
@@
 This is equivalent to FaviComp with $\alpha=0$, as we depend entirely on the compression model without any intervention from the target model.
-A detailed explanation of the implementation of the baselines is provided in Appendix A.
+[^3] A more detailed explanation of the implementation of the baselines is provided in Appendix A.
 Methods Size NQ TQA HQA Wiki MQ
@@
 indicates a fully-supervised compression model, where the compressor is trained.
-The best Accuracy and token-level F1 scores for each dataset are in bold.
 4 Experimental Results
-In this section, we compare the overall performance of FaviComp with other baselines across the five datasets (Section 4.1), explore the impact of ensemble coefficient $\alpha$ on performance and perplexity (Section 4.2), investigate how effectively FaviComp incorporate parametric and non-parametric knowledge (Section 4.3), and compare the compression rates with other baselines (Section 4.4).
+In this section, we compare the overall performance of FaviComp with other baselines across the five datasets (Section˜4.1), explore the impact of ensemble coefficient $\alpha$ on performance and perplexity (Section˜4.2), investigate how effectively FaviComp incorporate parametric and non-parametric knowledge (Section˜4.3), and compare the compression rates with other baselines (Section˜4.4).
 4.1 Main Results
-The overall performance of FaviComp and the baselines across the five datasets are presented in Table 1.222We present additional experimental results using other combinations of compression and target model at Section B.1.
+The overall performance of FaviComp and the baselines across the five datasets are presented in Table˜1.444We present additional experimental results using other combinations of compression and target model at Section B.1.
 To start with, the compression-based methods consistently outperform the reranking-based methods, due to the fact that the reranking-based methods are prone to losing more question-relevant information by discarding lower-ranked sentences.
@@
 Finally, given that Zero-shot Summarization corresponds to FaviComp with $\alpha=0$ and Generated Context corresponds to FaviComp with $\alpha=1$, the fact that FaviComp outperforms both baselines highlights its ability to effectively incorporate tokens from both sources—evidence summary and generated context.
-This results in superior performance compared to relying on just one source alone.
+This results in superior performance compared to relying on one source alone.
 Figure 2: Impact of coefficient $\alpha$ on performance and perplexity when using Llama3.2-3B-Instruct and Llama3-8B-Instruct compression-target pairs.
 4.2 Impact of Ensemble Coefficient on Performance and Perplexity
-Figure 2 illustrates how performance and perplexity change as the ensemble coefficient $\alpha$ is varied across the values when using Llama3.2-3B-Instruct and Llama3-8B- Instruct compression-target pairs on HQA and MQ datasets [^4] Results for other datasets are included in Figure 6..
+Figure˜2 illustrates how performance and perplexity change as the ensemble coefficient $\alpha$ is varied across the values when using Llama3.2-3B-Instruct and Llama3-8B- Instruct compression-target pairs on HQA and MQ datasets [^6] Results for other datasets are included in Figure 6..
 We calculate the perplexity of the compressed evidence conditioned on the preceding inputs, i.e.
@@
 Figure 3: Accuracy of baselines methods on $\mathrm{Hits=0}$ and $\mathrm{Hits=1}$ subset of multi-document QA datasets.
+Methods NQ TQA HQA Wiki MQ Generated Context 36.6 62.4 29.7 28.7 14.8 Zero-shot Summarization 43.2 67.1 38.5 31.1 15.3 Concatenation 42.5 66.7 36.5 29.0 15.6 FaviComp 46.8 70.9 41.6 35.2 19.9
+Table 2: Performance (F1) comparison against concatenation of parametric and non-parametric knowledge.
+4.3 Integration of Parametric and Non-parametric Knowledge
+The effective integration of parametric and non-parametric knowledge is crucial for complex tasks such as multi-document QA, where the evidence set may not contain all the necessary information.
+To this end, we evaluate how effectively FaviComp incorporates parametric knowledge from the target model and non-parametric knowledge from the compression model on the multi-document QA datasets.
+We begin by dividing the test samples of each dataset into evidence-relevant and evidence-irrelevant subsets, using the $\mathrm{Hits}$ metric.
+The $\mathrm{Hits}$ metric is set to 1 (evidence-relevant) if the retrieved evidence set contains the correct answer, and 0 (evidence-irrelevant) if it does not.
+We then assess the downstream performance of each subset.
+The underlying intuition is that if a method performs better on the evidence-relevant subset, it suggests that the method is more effectively utilizing the provided evidential knowledge.
+Conversely, if a method excels on the evidence-irrelevant subset, it indicates that the method is more effectively leveraging parametric knowledge without relying on potentially irrelevant evidence.
+As shown in Figure˜3, we compare the accuracy of FaviComp with Llama3.2-3B-Instruct and Llama3-8B-Instruct compression-target pairs on $\mathrm{Hits=0}$ and $\mathrm{Hits=1}$ subsets with the top-performing baselines, Zero-shot Summarization and CompAct [^7] We provide results of FaviComp on various alpha values in Section B.3.
+FaviComp outperforms other baselines in the $\mathrm{Hits=0}$ subset while performing comparably with others in the $\mathrm{Hits=1}$ subset.
+This proves that FaviComp effectively relies on parametric knowledge rather than evidential knowledge when faced with irrelevant evidence, while maintaining similar effectiveness in utilizing evidential knowledge when relevant evidence is present.
+In addition, we conduct another experiment to demonstrate FaviComp’s superior ability to synergize two sources of knowledge.
+We compare it against a straightforward approach that concatenates parametric and non-parametric knowledge as context for downstream generation.
+Specifically, we concatenate the compressed evidence from the Zero-shot Summarization with the generated context from the Generated Context and use this concatenated context for evaluation.
+The results, shown in Table˜2, reveal that simple concatenation underperforms compared to the Zero-shot Summarization baseline.
+This suggests that naively merging non-parametric and parametric knowledge in-context can be less effective than relying solely on non-parametric knowledge.
+In contrast, FaviComp effectively integrates both knowledge sources during compression, leveraging their synergy to achieve superior performance.
+4.4 Compression Rate Comparisons
+[…]
 Methods NQ TQA HQA Wiki MQ LongLLMlingua 1.87 1.84 1.83 1.83 1.83 RECOMP-abstractive 17.96 17.79 19.72 32.06 32.05 CompAct 8.85 8.92 9.45 10.71 8.96 Zero-shot Summarization 14.12 17.12 18.75 21.39 16.19 FaviComp 16.43 22.40 22.55 23.10 18.95
-Table 2: Compression rates of compression-based baselines. The highest compression rate is indicated in bold, while the second highest is underlined.
+Table 3: Compression rates of the baselines and FaviComp.
 Question: This film is an adaption of a Jacques Offenbach’s opera that was written by a Hungarian British screenwriter?
 Methods (Compressed) Evidence Prediction Perplexity
-Raw Document …(skip)… The Tales of Hoffmann is a 1951 British Technicolor film adaptation of Jacques Offenbach’s opera “The Tales of Hoffmann", written, produced and directed by the team of Michael Powell and Emeric Pressburger working under the umbrella of their production company, The Archers.
-Emeric Pressburger 12.429
+Raw Document The Tales of Hoffmann is a 1951 British Technicolor film adaptation of Jacques Offenbach’s opera “The Tales of Hoffmann", written, produced and directed by the team of Michael Powell and Emeric Pressburger … Emeric Pressburger 12.429
 Zero-shot Summarization The 1951 film “The Tales of Hoffmann" is an adaptation of Jacques Offenbach’s opera, written, produced, and directed by Michael Powell and Emeric Pressburger.
@@
 Raw Document The Chronicle of Philanthropy is a magazine that covers the nonprofit world.
-…(skip)… It was founded in 1988 by editor Phil Semas and then managing editor Stacy Palmer.
-…(skip)… Philanthropy (magazine) Philanthropy is a quarterly magazine published by the Philanthropy Roundtable.
-First published as a newsletter in 1987, "Philanthropy" became a glossy magazine in 1996.
+… It was founded in 1988 by editor Phil … First published as a newsletter in 1987, "Philanthropy" became a glossy magazine in 1996.
 Philanthropy 4.856
@@
 Tokens with no coloring represent those that are the $\arg\max$ of both models.
-4.3 Integration of Parametric and Non-parametric Knowledge
-The effective integration of parametric and non-parametric knowledge is crucial for complex tasks such as multi-document QA, where the evidence set may not contain all the necessary information.
-To this end, we evaluate how effectively FaviComp incorporates parametric knowledge from the target model and non-parametric knowledge from the compression model on the multi-document QA datasets.
-We begin by dividing the test samples of each dataset into evidence-relevant and evidence-irrelevant subsets, using the $\mathrm{Hits}$ metric.
-The $\mathrm{Hits}$ metric is set to 1 (evidence-relevant) if the retrieved evidence set contains the correct answer, and 0 (evidence-irrelevant) if it does not.
-We then assess the downstream performance of each subset.
-The underlying intuition is that if a method performs better on the evidence-relevant subset, it suggests that the method is more effectively utilizing the provided evidential knowledge.
-Conversely, if a method excels on the evidence-irrelevant subset, it indicates that the method is more effectively leveraging parametric knowledge without relying on potentially irrelevant evidence.
-As shown in Figure 3, we compare the accuracy of FaviComp with Llama3.2-3B-Instruct and Llama3-8B-Instruct compression-target pairs on $\mathrm{Hits=0}$ and $\mathrm{Hits=1}$ subsets with the top-performing baselines, Zero-shot Summarization and CompAct [^5] We provide results of FaviComp on vaious alpha values in Section B.3.
-FaviComp outperforms other baselines in the $\mathrm{Hits=0}$ subset while performing comparably with others in the $\mathrm{Hits=1}$ subset.
-This proves that FaviComp effectively relies on parametric knowledge rather than evidential knowledge when faced with irrelevant evidence, while maintaining similar effectiveness in utilizing evidential knowledge when relevant evidence is present.
-Interestingly, even though CompAct generally performs better on the $\mathrm{Hits=1}$ subset compared to Zero-shot Summarization, it underperforms relative to Zero-shot Summarization on the $\mathrm{Hits=0}$ subset.
-This suggests that the training may have been biased towards utilizing solely evidential knowledge, rather than effectively leveraging both sources in synergy.
-4.4 Compression Rate Comparisons
-[…]
 5 Case Study
-Table 3 presents two examples from HQA to illustrate how FaviComp effectively familiarizes evidence while seamlessly integrating both parametric and non-parametric knowledge during evidence compression.
+Table˜4 presents two examples from HQA to illustrate how FaviComp effectively familiarizes evidence while seamlessly integrating both parametric and non-parametric knowledge during evidence compression.
 We compare its output with Raw Document, which does not apply any compression, and Zero-shot Summarization.
@@
 (2024).
-Most recently, Jiang et al.
-(2023a) and Cao et al.
-(2024) conduct token-level or embedding-based compression to preserve only the query-relevant information using a trained compressor.
-Xu et al.
+Most recently, Xu et al.
 (2024) and Yoon et al.
 (2024) train a compression model to generate an abstractive summary of the documents by distilling knowledge from larger language models.
-While these methods are successful to some extent, they often achieve suboptimal performance because the compressed context may be unfamiliar to the LM used in the downstream task due to differences in the internal knowledge and prompt preferences between the compression and the target model.
+While these methods are successful to some extent, they often achieve suboptimal performance because of the discrepancy between the compression model and the target model, leading unfamiliarity of the context.
 In contrast, FaviComp proactively compresses the evidence pieces in a way to lower the target model’s perplexity using an ensemble decoding technique without any training, thereby improving the downstream performance.
 Parametric and Non-parametric Knowledge in RAG.
-While there have been studies on the phenomena of LM’s utilization of both parametric and non-parametric knowledge sources, there is a lack of research focused on effectively synergizing both sources.
+There has been a lack of research focused on effectively combining both sources.
 A few of these efforts introduce counterfactual augmentation (Longpre et al., 2021; Fang et al., 2024; Zhang et al., 2024) and causal intervention (Zhou et al., 2023; Wang et al., 2023a) to mitigate knowledge conflict, which, however, requires explicitly knowing the features of the input that causes such conflict.
@@
 7 Conclusion
-In this study, we introduce FaviComp, a training-free, decoding-time evidence compression method designed to enhance RAG performance by consolidating retrieved evidence set to be more familiar to the target model, while seamlessly integrating parametric knowledge.
+In this study, we introduce FaviComp, a training-free, inference-time evidence compression method designed to enhance RAG performance by consolidating retrieved evidence set to be more familiar to the target model, while seamlessly integrating parametric knowledge.
 Our extensive experiments validate the effectiveness of FaviComp on open-domain QA tasks, showing significant improvements over recent evidence compression baselines in multiple datasets.
-Additionally, FaviComp’s model-agnostic nature allows it to be effortlessly incorporated into various RAG workflows without additional training, making it a versatile tool for enhancing LMs in complex tasks.
+Additionally, FaviComp’s model-agnostic nature allows it to be incorporated into various RAG workflows at inference time, making it a versatile tool for enhancing LMs in complex tasks.
+Acknowledgment
+We appreciate the reviewers for their insightful comments and suggestions.
+This work was partly supported by the Amazon Nova Trusted AI Prize, the NSF of the United States Grants ITE 2333736 and OAC 2531126, and the DARPA FoundSci Grant HR00112490370.
 Limitations
@@
 However, it is a training-free strategy that can be easily plugged into any RAG application.
+We provide insights on the tradeoff between latency and performance in Section˜B.4.
 (2) Ensemble decoding requires the compression and target model to share the same vocabulary and tokenizer, which can limit the range of compatible models.
@@
 Appendix A Implementation Details
+A.1 Generation Configuration
+For all the baselines and FaviComp, we use default temperature and top-p values of the compression model during evidence compression and fix the temperature of the target model to 1.0 during evaluation.
+A.2 Dataset Statistics
+We provide the statistics of the evaluation dataset utilized in our experiments in Table˜6.
+A.3 Implementation Details of Baselines
 (1) Gold Compression: We implement the Gold Compression baseline following the approach outlined by Yoon et al.
@@
 This approach is necessary because the documents are chunked, and the retrieved documents may not exactly match the gold documents.
-(2) Generated Context: We use the context generation prompt in Table 7 to generate the context.
-(3) Zero-shot Summarization: We use the evidence compression prompt in Table 7 to compress the retrieved documents.
+(2) Generated Context: We use the context generation prompt in Table˜10 to generate the context.
+(3) Zero-shot Summarization: We use the evidence compression prompt in Table˜10 to compress the retrieved documents.
 (4) RECOMP-extractive: We utilize the same Contriever models trained by the authors for each dataset, to encode both the question and the sentences in the evidence set.
@@
 We conduct an experiment where we use Llama3 -8B-Instruct and Mistral-7B-Instruct for both compression and target models.
-The result in Table 5 demonstrates that FaviComp outperforms all other baselines, supplementing the effectiveness shown in Section 4.1.
+The result in Table˜8 demonstrates that FaviComp outperforms all other baselines, supplementing the effectiveness shown in Section˜4.1.
 B.2 Head-to-Head Comparison with RECOMP-abstractive
@@
 We train for 7 epochs using LoRA with Adam optimizer with a learning rate of 2e-6 and a batch size of 64.
-We present the evaluation results in Table 4.
+We present the evaluation results in Table˜5.
 Even though using larger base model for compression enhances the performance of RECOMP-abstractive to some extent, it still underperforms compared to training-free FaviComp.
@@
 We evaluate FaviComp’s performance on evidence-relevant ($\mathrm{Hits=1}$) and evidence-irrelevant ($\mathrm{Hits=0}$) subsets by varying $\alpha$ values.
-Figure 4 shows that $\alpha=0.5$ or $\alpha=0.7$ performs the best on the $\mathrm{Hits=0}$ subset, while performance declines as $\alpha$ deviates further from the value.
+Figure˜4 shows that $\alpha=0.5$ or $\alpha=0.7$ performs the best on the $\mathrm{Hits=0}$ subset, while performance declines as $\alpha$ deviates further from the value.
 This pattern in the $\mathrm{Hits=0}$ subset mirrors the overall performance trend, suggesting that appropriately utilizing parametric knowledge when the evidence is irrelevant is crucial to the overall performance.
@@
 Figure 4: Accuracy of FaviComp with various $\alpha$ values on $\mathrm{Hits=0}$ and $\mathrm{Hits=1}$ subset of multi-document QA datasets.
+Methods Compression Model Avg latency per sample (s) Performance
+Acc F1
+RECOMP-abstractive T5-large 0.22 39.3 43.3
+CompAct Mistral-7B-Instruct-v0.2 8.72 42.3 46.1
+Zero-shot Summarization Llama-3.2-3B-Instruct 3.99 39.4 43.2
+FaviComp Llama-3.2-3B-Instruct 6.43 42.8 46.8
+Table 7: Latency and of the baselines and FaviComp
+B.4 Latency Ablation Study
+Table˜7 shows the latency of our method along with other major baselines to provide insights on the trade-offs between accuracy and latency.
+We used Llama-3-8B-Instruct as the target model and tested on NQ dataset for the experiment.
+Although there are trade-offs between latency and accuracy across all methods, training-free FaviComp demonstrates lower latency while achieving higher accuracy than CompAct, which is the supervised baseline that previously achieved SOTA performance.
 Appendix C Prompt Templates
@@
 C.1 Evaluation
-The evaluation prompt template is shown in Figure 5.
+The evaluation prompt template is shown in Figure˜5.
 For all the evaluations throughout the experiment, we switch the positions of the Question and Context if doing so results in better performance.
-System prompts and demonstrations used in the evaluations are presented in Table 6 and Table 8, respectively.
+System prompts and demonstrations used in the evaluations are presented in Table˜9 and Table˜11, respectively.
 C.2 FaviComp
-The prompt templates for evidence compression and context generation of FaviComp are presented in Table 7
+The prompt templates for evidence compression and context generation of FaviComp are presented in Table˜10.
 Appendix D Licenses